\begin{thebibliography}{15}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Atlas et~al.(1990)Atlas, Cohn, and Ladner]{stream-based-sampling}
L.~Atlas, D.~Cohn, and R.~Ladner.
\newblock Training connectionist networks with queries and selective sampling.
\newblock In \emph{Advances in Neural Information Processing Systems 2}, pages
  566--573. Morgan Kaufmann Publishers Inc. San Francisco, CA, USA, 1990.

\bibitem[Brinker(2006)]{svm-binary-minimum}
K.~Brinker.
\newblock On active learning in multi-label classification.
\newblock In Myra Spiliopoulou, Rudolf Kruse, Christian Borgelt, Andreas
  N{\"u}rnberger, and Wolfgang Gaul, editors, \emph{From Data and Information
  Analysis to Knowledge Engineering}, pages 206--213, Berlin, Heidelberg, 2006.
  Springer Berlin Heidelberg.
\newblock ISBN 978-3-540-31314-4.

\bibitem[Buitinck et~al.(2013)Buitinck, Louppe, Blondel, Pedregosa, Mueller,
  Grisel, Niculae, Prettenhofer, Gramfort, Grobler, Layton, VanderPlas, Joly,
  Holt, and Varoquaux]{sklearn_api}
L.~Buitinck, G.~Louppe, M.~Blondel, F.~Pedregosa, A.~Mueller, O.~Grisel,
  V.~Niculae, P.~Prettenhofer, A.~Gramfort, J.~Grobler, R.~Layton,
  J.~VanderPlas, A.~Joly, B.~Holt, and G.~Varoquaux.
\newblock {API} design for machine learning software: experiences from the
  scikit-learn project.
\newblock In \emph{ECML PKDD Workshop: Languages for Data Mining and Machine
  Learning}, pages 108--122, 2013.

\bibitem[Cardoso et~al.(2017)Cardoso, Silva, Canuto, Moro, and
  Gonçalves]{ranked-batch-mode}
T.N.C. Cardoso, R.M. Silva, S.~Canuto, M.M. Moro, and M.A. Gonçalves.
\newblock Ranked batch-mode active learning.
\newblock \emph{Information Sciences}, 379:\penalty0 313--337, February 2017.

\bibitem[Cohn et~al.(1994)Cohn, Atlas, and Ladner]{query-by-disagreement}
D.~Cohn, L.~Atlas, and R.~Ladner.
\newblock Improving generalization with active learning.
\newblock \emph{Machine Learning}, 15\penalty0 (2):\penalty0 201--221, May
  1994.
\newblock ISSN 1573-0565.
\newblock \doi{10.1007/BF00993277}.
\newblock URL \url{https://doi.org/10.1007/BF00993277}.

\bibitem[Esuli and Sebastiani(2009)]{esuli-sebastiani}
A.~Esuli and F.~Sebastiani.
\newblock Active learning strategies for multi-label text classification.
\newblock In Mohand Boughanem, Catherine Berrut, Josiane Mothe, and Chantal
  Soule-Dupuy, editors, \emph{Advances in Information Retrieval}, pages
  102--113, Berlin, Heidelberg, 2009. Springer Berlin Heidelberg.
\newblock ISBN 978-3-642-00958-7.

\bibitem[Lewis and Catlett(1994)]{uncertainty-sampling}
D.~D. Lewis and J.~Catlett.
\newblock Heterogeneous uncertainty sampling for supervised learning, 1994.

\bibitem[Li et~al.(2004)Li, Wang, and Sung]{svm-max-loss}
X.~Li, L.~Wang, and E.~Sung.
\newblock Multilabel svm active learning for image classification.
\newblock In \emph{2004 International Conference on Image Processing, 2004.
  ICIP '04.}, volume~4, pages 2207--2210 Vol. 4, Oct 2004.
\newblock \doi{10.1109/ICIP.2004.1421535}.

\bibitem[McCallum and Nigam(1998)]{information-density}
A.~McCallum and K.~Nigam.
\newblock Employing em and pool-based active learning for text classification.
\newblock In \emph{'98 Proceedings of the Fifteenth International Conference on
  Machine Learning (ICML)}, pages 350--358. Morgan Kaufmann Publishers Inc. San
  Francisco, CA, USA, 1998.

\bibitem[Pedregosa et~al.(2011)Pedregosa, Varoquaux, Gramfort, Michel, Thirion,
  Grisel, Blondel, Prettenhofer, Weiss, Dubourg, Vanderplas, Passos,
  Cournapeau, Brucher, Perrot, and Duchesnay]{scikit-learn}
F.~Pedregosa, G.~Varoquaux, A.~Gramfort, V.~Michel, B.~Thirion, O.~Grisel,
  M.~Blondel, P.~Prettenhofer, R.~Weiss, V.~Dubourg, J.~Vanderplas, A.~Passos,
  D.~Cournapeau, M.~Brucher, M.~Perrot, and E.~Duchesnay.
\newblock Scikit-learn: Machine learning in {P}ython.
\newblock \emph{Journal of Machine Learning Research}, 12:\penalty0 2825--2830,
  2011.

\bibitem[Reyes et~al.(2016)Reyes, P{{\'e}}rez, del Carmen
  Rodr{{\'i}}guez-Hern{{\'a}}ndez, Fardoun, and Ventura]{JCLAL}
O.~Reyes, E.~P{{\'e}}rez, M~del Carmen Rodr{{\'i}}guez-Hern{{\'a}}ndez, H.~M.
  Fardoun, and Sebasti{\'a}n Ventura.
\newblock {JCLAL}: A java framework for active learning.
\newblock \emph{Journal of Machine Learning Research}, 17\penalty0
  (95):\penalty0 1--5, 2016.
\newblock URL \url{http://jmlr.org/papers/v17/15-347.html}.

\bibitem[Roy and McCallum(2001)]{expected-error}
N.~Roy and A.~McCallum.
\newblock Toward optimal active learning through sampling estimation of error
  reduction.
\newblock In \emph{Proceedings of the International Conference on Machine
  Learning (ICML)}, pages 441--448. Morgan Kaufmann Publishers Inc. San
  Francisco, CA, USA, 2001.

\bibitem[Seung et~al.(1992)Seung, Oper, and Sompolinsky]{query-by-committee}
H.S. Seung, M.~Oper, and H.~Sompolinsky.
\newblock Query by committee.
\newblock In \emph{Proceedings of the ACM Workshop on Computational Learning
  Theory}, pages 287--294, 1992.
\newblock \doi{10.1145/130385.130417}.

\bibitem[Shahriari et~al.(2016)Shahriari, Swersky, Wang, Adams, and
  de~Freitas]{7352306}
B.~Shahriari, K.~Swersky, Z.~Wang, R.~P. Adams, and N.~de~Freitas.
\newblock Taking the human out of the loop: A review of {B}ayesian
  optimization.
\newblock \emph{Proceedings of the IEEE}, 104\penalty0 (1):\penalty0 148--175,
  Jan 2016.
\newblock ISSN 0018-9219.
\newblock \doi{10.1109/JPROC.2015.2494218}.

\bibitem[Yang et~al.(2017)Yang, Lee, Chung, Wu, Chen, and Lin]{YY2017}
Y.-Y. Yang, S.-C. Lee, Y.-A. Chung, T.-E. Wu, S.-A. Chen, and H.-T. Lin.
\newblock libact: Pool-based active learning in {P}ython.
\newblock Technical report, 2017.
\newblock arXiv: \url{https://arxiv.org/abs/1710.00379}.

\end{thebibliography}

\documentclass[twoside,11pt]{article}

\usepackage{jmlr2e}
\usepackage{graphicx}
\usepackage{float}
\usepackage{booktabs}

\usepackage{listings}
\usepackage{color}

\definecolor{dkgreen}{rgb}{0,0.6,0}
\definecolor{gray}{rgb}{0.5,0.5,0.5}
\definecolor{mauve}{rgb}{0.58,0,0.82}

\graphicspath{{fig/}}

\ShortHeadings{modAL: A modular active learning framework for Python}{}
\firstpageno{1}

\begin{document}

\title{modAL: A modular active learning framework for Python}

\author{\name Tivadar Danka \email danka.tivadar@brc.mta.hu \\
       \addr Biological Research Centre\\
       Hungarian Academy of Sciences\\
       Szeged, H6720, Hungary
       \AND
       \name Peter Horvath \email horvath.peter@brc.mta.hu \\
       \addr Biological Research Centre\\
	   Hungarian Academy of Sciences\\
	   Szeged, H6720, Hungary}
	   
\maketitle

\begin{abstract}
modAL is a modular active learning framework for Python, aimed to make active learning research and practice simpler. Its distinguishing features are (i) clear and modular object oriented design (ii) full compatibility with scikit-learn models and workflows. These features make fast prototyping and easy extensibility possible, aiding the development of real-life active learning pipelines and novel algorithms as well. modAL is fully open source, hosted on GitHub.\hspace{1pt}\footnote{https://github.com/modAL-python/modAL} To assure code quality, extensive unit tests are provided and continuous integration is applied. In addition, a detailed documentation with several tutorials are also available for ease of use. The framework is available in PyPI and distributed under the MIT license.
\end{abstract}

\begin{keywords}
	Active Learning, scikit-learn, Machine Learning, Python
\end{keywords}

\section{Introduction}
Upon learning patterns from data in real-life applications, labelling examples often consume significant time and money, which makes it infeasible to obtain large training sets. For example, sentiment analysis of texts requires extensive manual annotation, which costs expert time. Another example is the optimization of black box functions, for which the evaluation is costly or derivatives are not available. In these cases, active learning can be used to query labels for the most informative instances. modAL is an active learning framework for Python, designed with modularity, flexibility and extensibility in mind. Built on top of scikit-learn (\cite{scikit-learn, sklearn_api}), it allows the rapid prototyping of active learning workflows with a large degree of freedom. It was designed to be easily extensible, allowing researchers to implement and test novel active learning strategies with minimal effort.

\section{Design principles and features}
Our objective with modAL was to create an active learning library which takes advantage of the advanced features of Python and the extensive ecosystem of scikit-learn, making the implementation of complex workflows simple and intuitive. Specifically, modAL was designed with the following goals in mind. \\
\begin{enumerate}
	\item \textbf{Modularity: separating and recombining parts of a workflow.} In general, an active learning workflow consists of a learning algorithm and a query strategy. In modAL, this is represented by the \verb|ActiveLearner| class, for which these components are passed upon object creation. Learning algorithms can be used with query strategies in any combination, making rapid prototyping possible.	
	\item \textbf{Extensibility: simple customization of parts.} In a modAL active learning workflow, a query strategy is simply a function, given to the object representing the active learning algorithm upon initialization. Implementing custom query strategies can be done without understanding class structures or modAL internals. Thus it requires minimal effort, allowing researchers to easily test novel strategies and compare them with existing ones. 
	\item \textbf{Flexibility: compatibility with the scikit-learn ecosystem.} scikit-learn is one of the most popular machine learning tools in Python, used by researchers and practicioners as well. modAL is built on top of it, allowing the use any of its classifier and regressor algorithms in active learning pipelines. Objects in modAL also follow the scikit-learn API, making it possible to insert them into already existing workflows.
\end{enumerate}

modAL supports a wide range of active learning algorithms for both pool-based and stream-based (\cite{stream-based-sampling}) scenarios. For multiclass problems, uncertainty sampling methods such as least confident (\cite{uncertainty-sampling}), max margin and max entropy sampling; committee-based methods such as query by committee (\cite{query-by-committee}) and query by disagreement (\cite{query-by-disagreement}); ranked batch-mode sampling (\cite{ranked-batch-mode}); expected error reduction (\cite{expected-error}) is provided. For multilabel classification, SVM binary minimum (\cite{svm-binary-minimum}); max loss and mean max loss (\cite{svm-max-loss}); MinConfidence, AvgConfidence, MinScore, AvgScore (\cite{esuli-sebastiani}) algorithms are implemented. For density weighting, the information density framework (\cite{information-density}) is available. In addition to classification, active regression. Moreover, Bayesian optimization is available with probability of improvement, expected improvement and upper confidence bound strategies (\cite{7352306}). 

\section{Classes and interfaces}\label{section:classes-and-interfaces}

For modularity and easy extensibility, active learning workflows are abstracted and represented by the \verb|ActiveLearner|, \verb|BayesianOptimizer|, \verb|Committee| and \verb|CommitteeRegressor| classes. All classes inherit from the \verb|sklearn.base.BaseEstimator| class. \verb|ActiveLearner| serves as an abstract model for general active learning algorithms, while \verb|Committee| and \verb|CommitteeRegressor| implements committee-based strategies. Bayesian optimization algorithms are represented by \verb|BayesianOptimizer|. All classes require a learner and a query strategy upon initialization. In the case of \verb|ActiveLearner| and \verb|BayesianOptimizer|, the learner is an arbitrary object implementing the scikit-learn API, while the for \verb|Committee| and \verb|CommitteeRegressor|, a list of \verb|ActiveLearner| instances must be provided. Again, the query function can be factored into two functions: one calculating the utility for each instance and one selecting the instances to be queried based upon the utility score. This modular design makes easy extensibility and interaction with other libraries possible. The use of ActiveLearner is demonstrated below. \\

\begin{lstlisting}
from modAL.models import ActiveLearner
from modAL.uncertainty import uncertainty_sampling
from sklearn.ensemble import RandomForestClassifier

# initializing the learner
learner = ActiveLearner(
    estimator=RandomForestClassifier(),
    query_strategy=uncertainty_sampling
)

# training
learner.fit(X_training, y_training)

# query for labels
query_idx, query_inst = learner.query(X_pool)
# ...obtaining new labels from the Oracle...
# supply label for queried instance
learner.teach(X_pool[query_idx], y_new)
\end{lstlisting}


\section{Comparison with other libraries}
To assess the features of modAL, a comparison between libraries is provided. We compare modAL to the Python libraries acton\hspace{1pt}\footnote{https://github.com/chengsoonong/acton}, alp\hspace{1pt}\footnote{https://github.com/davefernig/alp}, libact\hspace{1pt}\footnote{https://github.com/ntucllab/libact} (\cite{YY2017}) and the Java library JCLAL (\cite{JCLAL}) in Tables \ref{tab:algorithms_full}, \ref{tab:other-features}. The comparison is made with respect to supported algorithms, design and support. For Python libraries, a runtime comparison for least confident sampling, query by committee\footnote{A minor bugfix was applied on \textit{alp} for QBC to work, see https://github.com/davefernig/alp/issues/1} and expected error reduction can be found in Table \ref{tab:runtime}. The runtime data was obtained by averaging the result of 10 runs for each algorithm. During each run, 10 queries were made. The comparison script is available at https://github.com/modAL-python/modAL/blob/master/examples/runtime\_comparison.py.

\section{Availability}
The framework is fully open-source, hosted on GitHub.\hspace{1pt}\footnote{https://github.com/modAL-python/modAL} Besides the core features, detailed documentation and a wealth of examples and tutorials are available at the project website\hspace{1pt}\footnote{https://modAL-python.github.io}, making active learning accessible for a wide range of users. All tutorials and examples on the official website can be downloaded as a Jupyter notebook. To assure code quality, extensive unit tests are provided with $ 98\% $ code coverage. Continuous integration is applied using Travis-CI. modAL is also available from PyPI.

\acks{T.D. and P.H. acknowledges support from the European Regional Development Funds (GINOP-2.3.2-15-2016-00001, GINOP-2.3.2-15-2016-00037).}

\begin{table}\centering
	\begin{tabular}{r c c c c c c c c c c c c c c c} \toprule
		& \rotatebox[origin=c]{90}{pool} 
		& \rotatebox[origin=c]{90}{stream} 
		& \rotatebox[origin=c]{90}{regression} 
		& \rotatebox[origin=c]{90}{committee} 
		& \rotatebox[origin=c]{90}{multilabel} 
		& \rotatebox[origin=c]{90}{information} \rotatebox[origin=c]{90}{density} 
		& \rotatebox[origin=c]{90}{expected} \rotatebox[origin=c]{90}{error} 
		& \rotatebox[origin=c]{90}{variance} \rotatebox[origin=c]{90}{reduction} 
		& \rotatebox[origin=c]{90}{hierarchical} \rotatebox[origin=c]{90}{sampling} 
		& \rotatebox[origin=c]{90}{ranked batch} 
		& \rotatebox[origin=c]{90}{meta-} \rotatebox[origin=c]{90}{learning} 
		& \rotatebox[origin=c]{90}{multilabel} 
		& \rotatebox[origin=c]{90}{Bayesian} \rotatebox[origin=c]{90}{optimization}
		& \rotatebox[origin=c]{90}{cost} \rotatebox[origin=c]{90}{sensitive} \\ \midrule
		modAL & \checkmark & \checkmark & \checkmark & \checkmark & \checkmark & \checkmark & \checkmark & X & X & \checkmark & X & \checkmark & \checkmark & X\\
		acton & \checkmark & X & \checkmark & \checkmark & X & X & X & X & X & X & X & X & X & X \\
		alp & \checkmark & X & X & \checkmark & X & X & X & X & X & X & X & X & X & X \\
		libact & \checkmark & X & X & \checkmark & \checkmark & X & \checkmark & \checkmark & \checkmark & X & \checkmark & \checkmark & X & \checkmark \\
		JCLAL & \checkmark & \checkmark & X & \checkmark & \checkmark & \checkmark & \checkmark & \checkmark & X & X & X & \checkmark & X & X \\ \bottomrule
	\end{tabular}
	\caption{Comparison of libraries with respect to supported algorithms.}
	\label{tab:algorithms_full}
	
	\vspace{2em}
	
%
	
	\begin{tabular}{r c c c c c} \toprule
		& \vtop{\hbox{\strut sklearn}\hbox{\strut model usability}} & \vtop{\hbox{\strut follows}\hbox{\strut sklearn API}} & \vtop{\hbox{\strut actively}\hbox{\strut maintained}} & \vtop{\hbox{\strut Python}\hbox{\strut version}} & \vtop{\hbox{\strut documentation,}\hbox{\strut tutorials}}\\ \midrule
		modAL & \checkmark & \checkmark & \checkmark & 3 & \checkmark \\ 
		acton & with adapters & \checkmark & \checkmark & 3 & \checkmark \\
		alp & \checkmark & X & X & 2, 3 & X \\
		libact & with adapters & X & \checkmark & 2, 3 & \checkmark \\
		JCLAL & -- & -- & \checkmark & -- & \checkmark \\
		\bottomrule
	\end{tabular}
	\caption{Comparison of libraries with respect to design and support.}
	\label{tab:other-features}
	
	\vspace{2em}
	
	\begin{tabular}{r c c c} \toprule
		& least confident & QBC & EER \\ \midrule
		modAL & 0.0087 s & 0.0465 s & 2.1255 s \\
		acton & 0.1860 s & 0.5858 s & -- \\
		alp & 0.0055 s & 0.0573 s & -- \\
		libact & 0.0191 s & 0.0324 s & 2.8436 s \\ \bottomrule
	\end{tabular}
	\caption{Comparison of Python libraries with respect to runtime. For each algorithm, 10 queries were made in a run and each run was repeated 10 times.}
	\label{tab:runtime}
\end{table}


\vskip 0.2in
\bibliography{modAL}

\end{document}